OpenStudy (anonymous):

You are given the parametric equations of a curve and a value for the parameter t. Find the coordinates of the point on the curve corresponding to the given value of t. x = 9 sin3 t, y = 9 cos3 t; t = π/4 (x, y)= ?

OpenStudy (campbell_st):

well you make the substitution to find x and y \[x = 9 \sin(3 \times \frac{\pi}{4})~~~~y = 9 \times \cos(3 \times \frac{\pi}{4})\] both trig ratios will have exact values and are both in the 2nd quadrant.. so one is positive and 1 is negative.